OpenStudy (beth12345):

How to find 'X' when sin90/3.6 = sinX/2.0 ?

OpenStudy (anonymous):

is that 90 as in 90 degrees?

OpenStudy (beth12345):

yup

OpenStudy (anonymous):

\[\sin(90)=1\] so you have \[\frac{1}{3.6}=\frac{\sin(x)}{2}\] \[\sin(x)=\frac{2}{3.6}\] \[x=\sin^{-1}(\frac{2}{3.6})\] then a calculator

OpenStudy (beth12345):

I get 0.589

OpenStudy (beth12345):

the answer in the back is 34 but I just don't know how to get it

OpenStudy (anonymous):

the ans is 34 when u round up, r u sure u typed into the calculator correctly?

OpenStudy (beth12345):

uhm i don't know ill try again

OpenStudy (beth12345):

i got the same thing, i think mabe my calculator is on a wrong setting or something, ill try to fix it

OpenStudy (beth12345):

oh ya i fixed it now, i got 33.749 and rounded up = 34

OpenStudy (anonymous):

that's good! :)

OpenStudy (beth12345):

thanks :)

OpenStudy (anonymous):

no prob :P